OpenStudy (mackenzie2013):

In Chad's reading class, all the students are reading the same book. The school bought each student a book at $7 per book. If there are 27 students in Chad's class, which of the following expressions could not be used to calculate the total cost? 7(30) - 7(7) 7(20) + 7(7) 7(30) - 7(3)

hartnn (hartnn):

7*27 = 7 (30-3) = 7(30) = 7(3)